Fill in the blanks with the missing numbers.
1. 5(x²-2x+___)=0+__
2. 2(x²-6x+__)=0+__

Answer by fractalier(6550)

this looks like an exercise on completing the square...if so, we have
1. 5(x²-2x+ 1)=0+ 5 (half of -2 squared = 1
2. 2(x²-6x+ 9)=0+ 18 (half of -6 squared = 9)